### v1

@@ Abstract
Aims. We introduce an estimation tool within the Gammapy landscape, an open-source Python package for the analysis of gamma-ray data, to quantify the energy-dependent morphology of a gamma-ray source.
Methods. The proposed method for this estimation tool fits the spatial morphology in a global fit across all energy slices (null hypothesis) and compares this to separate fits for each energy slice (alternative hypothesis). These are modelled using forward-folding methods, and the significance of variability is quantified by comparing the test statistic between the two hypotheses.
Results. We present a general tool to probe changes in the spatial morphology with energy, employing a full forward folding approach with 3D likelihood. We present its usage on a real dataset from H.E.S.S., and also on a simulated dataset to quantify the significance of energy dependence for sources of different sizes. In the first example, which utilises a subset of data from HESS J1825--137, we observe extended emission at lower energies that becomes more compact at higher energies. The tool indicates very significant variability ($9.8\sigma$) in the case of the largely extended emission. In the second example, a source with a smaller extent ($\sim 0.1^{\circ}$), simulated using the CTAO response, shows the tool can still provide a statistically significant variation ($9.7\sigma$), despite the small scale.
@@ 1 Introduction
Detecting variations in the morphology of gamma-ray sources with energy can offer useful insight into their underlying physical mechanisms, however, measuring and quantifying these variations remains difficult. These energy-dependent morphologies are expected in a variety of astrophysical sources, including pulsar wind nebulae (PWNe), supernova remnants (SNRs) and relativistic jets. It can therefore be important to look for any evidence of this energy dependence.
@@
The evolutionary phases of a PWN lead to a unique gamma-ray morphology (Gaensler & Slane, 2006). The emission is expected to be compact and close to the pulsar position at high energies, however, for lower energies the emission is expected to be more extended with an offset from the pulsar position. Radiative cooling, either by synchrotron emission in magnetic fields or inverse Compton scatter of ambient photon fields, leads to higher energy electrons losing energy faster. This creates a spectral steepening with distance. Another physical process at play is the energy-dependent diffusion of particles, in which high energy particles diffuse more efficiently than the lower energy ones. Energy-dependent morphology can also be attributed to the bulk motion of particles in the PWN outflow. Advection dominated transport in this case, leads to a spatial distribution dependent on the energy loss timescale. Conversely, for SNRs with uniform gas densities, we expect the gamma-ray emission at lower energies to be concentrated to the central SNR position, and the high-energy emission to be further away (Celli et al., 2019; Brose et al., 2021). Relativistic jets, such as those observed in microquasars, are also expected to exhibit energy-dependent morphology along their extent (H. E. S. S. Collaboration et al., 2024).
@@
The High Energy Stereoscopic System (H.E.S.S.) has detected several sources exhibiting energy-dependent morphologies. Examples include evolved PWNe such as HESS J1303--631 (H. E. S. S. Collaboration et al., 2012a), and HESS J1825--137, where the nebula size increases with decreasing energy. HESS J1825--137 is a well-known example of a PWN that shows a clear energy-dependent gamma-ray morphology (Aharonian et al., 2006a; H.E.S.S. Collaboration et al., 2019; Principe et al., 2020). A compact nebula with an extended diffuse nebula was discovered in X-rays towards PSR B1823--13 with asymmetric morphology (Finley et al., 1996). Aharonian et al. (2005) discovered an extended very-high-energy (VHE) gamma-ray nebula towards this region, designated HESS J1825--137. The size of the nebula is seen to increase with decreasing energy (H.E.S.S. Collaboration et al., 2019). At low energies ($<32$ TeV) the emission nebula is extended with its peak offset from the position of PSR B1823--13. For higher energies ($>32$ TeV), the emission is more compact and close to the pulsar position. This behaviour is indicative that the gamma-ray emission towards HESS J1825--137 originates from PSR B1823--13 as a PWN.
@@
SS 433 is the first microquasar to be detected at TeV energies (Abeysekara et al., 2018; H. E. S. S. Collaboration et al., 2024). The H.E.S.S. gamma-ray image, with the nearby extended emission from HESS J1908$+$063 modelled out, reveals two regions of emission, known as the eastern and western jets. No TeV emission is significantly detected from the central source. At the highest energies (¿10 TeV) the emission appears only in the base of the X-ray jets, i.e. the point closest to the central source. In contrast, the lower energy gamma-rays (0.8–2.5 TeV and 2.5–10 TeV) peak in significance at positions further along the jets.
@@
Various techniques, applicable in different scenarios, have been utilised to search for energy dependence. These include the comparison of spectral indices across different regions or energy bins in a map (Aharonian et al., 2006a), slicing spatial maps into different energy bands and comparing their extension from a model fit (see, for instance, examples of Gaussian spatial model fits in H. E. S. S. Collaboration et al., 2012a; Principe et al., 2020; MAGIC Collaboration et al., 2020), comparing radial profiles (Principe et al., 2020) and many other techniques. Energy-dependent morphology is a key observable to study the nature of a VHE gamma-ray source. Here, we present a new tool that was introduced in the Gammapy framework (Donath et al., 2023; Acero et al., 2024), to investigate the potential energy-dependent morphology towards a gamma-ray source. This method utilises maps across multiple energy bands to fit the morphological parameters of the source, such as position and extension. The user defines the best-fit model for the morphology and spectrum. A global fitting approach is applied, then an individual fit is performed in each energy band to assess if the morphology changes significantly.
@@ 2 Methodology
In this section, we describe various methods for measuring the energy dependence of gamma-ray sources, highlighting their respective advantages and disadvantages.
@@
One common approach is spatially resolved spectral analysis (Aharonian et al., 2006b, c). This technique can be used to search for a softening of the gamma-ray emission away from the pulsar position in a PWN scenario, which provides the evidence of an energy-dependent morphology. The signal region is divided into non-overlapping subregions. A spectral model, typically a power law, is fit in each subregion independently to search for changing photon index across the source region (Aharonian et al., 2006c). This approach offers a simple method of analysing the spectra, particularly for variations in spectral shape. However, this method has its limitations. The choice of region size and number of regions can strongly influence the result, leading to potential biases. The astrophysical background is difficult to determine accurately for each region, the estimation for different regions are typically not independent. These limitations are more significant for small subregions which approach the resolution of the point spread function (PSF), making this method better suited to more extended sources. Quantifying the significance of spectral variability across these subregions remains a challenge.
@@
Another method involves creating radial excess profiles along a specific axis (H. E. S. S. Collaboration et al., 2012b; H.E.S.S. Collaboration et al., 2019; Principe et al., 2020). The distribution of the emission is analysed by extracting the excess counts from specific regions in a specific orientation from the source position. This is useful to visualise changing spatial features such as extension and intensity. It is important to note that they do not take into account the PSF or variations in exposure. Radial profiles offer a model-independent method to assess the spatial variations, allowing for an unbiased way of visualising the data. However, they assume radial symmetry within the region, which may not always be a valid approximation. Additionally, this method is highly sensitive to the selected axis direction of the profile and the choice of the radial bins. Another limitation is the difficulty in handling background systematics. Consequently, these dependencies make it a less robust analysis. Additionally, variations in the exposure across the field of view can change the excess distribution, meaning the excess partially reflects these exposure variations rather than the source properties itself. For example, a decrease in excess counts may be misinterpreted as a physical or temporal effect, when it is, in fact, stemming from the exposure variation. This makes deriving the true significance of the effect a challenge. While traditionally overlooked, accounting for exposure variation is now possible with the introduction of methods like the FluxProfileEstimator in Gammapy. This, however, still does not account for contributions from additional sources within the field of view or analysing variations on small spatial scales.
@@
We propose a new method, that addresses each of the limitations described above, to quantify the energy-dependent morphology with a proper statistical assessment. This technique utilises a full forward-folding approach, where a morphological fit is conducted across various energy slices and compared with a global morphology fit. The technique chosen here involves fitting a spatial model and comparing the differences in morphological parameters (such as extension and position) in different energy bins. Further information about the tool and examples can be found in section 3 and section 4, respectively.
In gamma-ray astronomy we often utilise hypothesis testing. We rely on Wilks theorem (Wilks, 1938) which is a log-likelihood ratio test which allows us to estimate the relative significance through the test statistic:
@@
where $\mathcal{L}_{0}^{\rm{max}}$ and $\mathcal{L}_{1}^{\rm{max}}$ represent the maximum likelihood of the model under the null and alternative hypotheses, respectively. Under the correct conditions, where the two models are nested, and the regularity conditions are satisfied, we expect the distribution of $\Delta$TS to asymptotically approach a $\chi^{2}$ function with $m$ degrees of freedom, the difference in number of free parameters between the two nested hypotheses. Therefore, for a large enough sample it is possible to compute the $\Delta$TS and compare to a chi-squared function with the correct degrees of freedom to test the statistical significance.
For our tool, the null hypothesis (H0) is defined by a global fit. Initially, the morphological parameters are fit in the first energy band. The resulting parameters are then applied to all subsequent energy bands and a fit is performed across the different energy bands. The alternative hypothesis (H1) is defined by individual fits. Here, the dataset is divided into energy bins, and both the spatial and spectral models are fitted independently in each energy band. We apply Wilks theorem to these hypotheses to obtain the test statistic ($\Delta$TS, see Equation 1). The test statistic is converted into a significance value through the survival function of the chi-squared distribution. Thus, this method provides a quantitative measure of how much the morphology deviates from the null model for each energy band.
@@ 3 Implementation
To test for energy-dependent morphology using our tool we begin by dividing the MapDataset into different energy bins, as specified by the user. Firstly, we allow (at a minimum) the position and extension of the SpatialModel to be free, along with the amplitude of the SpectralModel and normalisation of the background. It is up to the user to define which spatial parameters are to be kept free. For example, for a GaussianSpatialModel there are five free parameters - longitude, latitude, sigma, phi, and e. It is up to the user to make a judicious choice to freeze/free these parameters. In this case, longitude, latitude, sigma are kept free. For a proper energy dependent investigation, it is necessary to keep at least the position and extension free. This enables us to detect any change in the position or size of the gamma-ray source across different energy bands. The spectral index can be poorly constrained, particularly when fitting in small energy bands. Therefore, it is recommended to assume a given spectral shape while allowing the amplitude to be fit. This re-normalisation can account for any small variations in the spectrum.
Whilst there are several SpatialModels available in Gammapy, it is necessary to select a model that accurately represents the morphology of the gamma-ray source. SpatialModels such as the DiskSpatialModel and ShellSpatialModel, assume sharp edges, which are often not a good representation of the data, particularly across multiple energy bands. In contrast, the GaussianSpatialModel is able to represent the smooth features observed in gamma-ray morphology, particularly when the morphology varies with energy. This makes it a robust default for many analyses. Alternative models should be used with caution and only when clearly justified by the data. A usage example is shown in Figure 1, corresponding to the example in subsection 4.2.
@@ 4.1 Example 1: HESS J1825--137
To demonstrate the functionality of the tool we showcase two examples. In the first example, we utilise a subset of the H.E.S.S. data towards HESS J1825--137 from Aharonian et al. (2006b). Three energy bins are chosen for this study: 0.4-2 TeV, 2-10 TeV and 10-100 TeV. For this specific example, a mask is applied to exclude the emission from the microquasar LS 5039 and the source HESS J1826--130. As shown by Figure 2, the dataset shows three distinct energy bins.
@@
Figure 2: Significance map of the H.E.S.S. dataset towards HESS J1825--137 in the various energy bins of interest. An oversampling radius of $0.1^{\circ}$ is used. The white circles show the mask applied to the emission microquasar LS 5039 and the gamma-ray source HESS J1826--130. The turquoise dot indicates the position of the pulsar, PSR B1823--13. We can see by-eye there is a clear change in the size of the emission in the different bins.
@@
Both a PowerLawSpectralModel and GaussianSpatialModel provide reasonable initial spectral and spatial models, respectively, for the dataset, so we therefore proceed with these for our study. The power law spectral shape is defined as follows:
$\phi(E)=\phi_{0}\cdot\left(\frac{E}{E_{0}}\right)^{-\Gamma}$ (2)
@@
where $\phi_{0}$ is the amplitude, $\Gamma$ is the index of the power law fit and $E_{0}$ is a normalisation energy taken here to be 1 TeV. We utilise the GaussianSpatialModel in Gammapy, keeping the asymmetry frozen to 0 for simplicity, in the different energy bands, to test for any energy-dependent morphology with our tool.
@@
An important aspect of this new tool is to check that the energy bands chosen contain significant data. To do this, we can look at the signal above the background in each energy bin. This is done by slicing the dataset into its energy bins. The null hypothesis (H0) involves freezing all parameters except for the normalisation of the background, which is fitted in each energy band. The alternative hypothesis (H1) allows additional parameters to be free, in this case the position and extension of the spatial model and the amplitude of the spectral model. Subsequently, the fit is performed for each energy band. These hypotheses are compared using Wilks theorem (Equation 1) to obtain a test statistic that indicates the significance of the signal relative to the background in each energy band. It was found that each energy bin is significant above the background ($\gg 5\sigma$), therefore we can continue with the energy-dependent study.
@@
The next step is to quantify any energy-dependent morphology. The null hypothesis is determined by applying the same set of parameters to all energy bands and performing the fit. In the alternative hypothesis, the free parameters of the model are fit individually within each energy band. We can plot the fitted spatial models from our energy dependent study. These are shown for each energy band in Figure 3. Each circle shows the extension and error of the fitted GaussianSpatialModels. The plus signs indicate the position and position error of the GaussianSpatialModel for each energy band.
@@
Figure 3: The results of the estimator tool for the HESS J1825--137 dataset. The different coloured rings show the fitted GaussianSpatialModel in the various energy bins along with their uncertainties. The plus symbols show the position and error from the fit. The black dot indicates the position of the pulsar, PSR B1823--13.
@@
Aharonian et al. (2006b) investigate the spectral properties of HESS J1825--137 in different regions to search for any energy-dependent morphology. A fit to these results gives a chi-squared value of 58 with 11 degrees of freedom. This shows that the method outlined in this paper is more sensitive.

### v3

@@ Abstract
Aims. We introduce an estimation tool within the Gammapy landscape, an open-source Python package for the analysis of gamma-ray data, for quantifying the energy-dependent morphology of a gamma-ray source.
Methods. The proposed method fits the spatial morphology in a global fit across all energy slices (null hypothesis) and compares this to separate fits for each energy slice (alternative hypothesis). These are modelled using forward-folding methods, and the significance of the variability is quantified by comparing the test statistics of the two hypotheses.
Results. We present a general tool for probing changes in the spatial morphology with energy, employing a full forward-folding approach with a 3D likelihood. We present its usage on a real dataset from H.E.S.S. and on a simulated dataset to quantify the significance of the energy dependence for sources of different sizes. In the first example, which utilises a subset of data from HESS J1825--137, we observe extended emission at lower energies that becomes more compact at higher energies. The tool indicates a very significant variability ($9.8\sigma$) in the case of the largely extended emission. In the second example, a source with a smaller extent ($\sim 0.1^{\circ}$), simulated using the CTAO response, shows the tool can still provide a statistically significant variation ($9.7\sigma$) on small scales.
@@ 1 Introduction
Detecting variations in the morphology of gamma-ray sources with energy can offer useful insight into their underlying physical mechanisms. However, measuring and quantifying these variations remains difficult. These energy-dependent morphologies are expected in a variety of astrophysical sources, including pulsar wind nebulae (PWNe), supernova remnants (SNRs), and relativistic jets. It can therefore be helpful to look for any evidence of this energy dependence.
@@
The evolutionary phases of a PWN lead to a unique gamma-ray morphology (Gaensler & Slane, 2006). The emission is expected to be compact and close to the pulsar position at high energies. However, for lower energies, the emission is expected to be more extended, with an offset from the pulsar position. Radiative cooling, either via synchrotron emission in magnetic fields or the inverse Compton scatter of ambient photon fields, leads to higher energy electrons losing energy faster. This creates a spectral steepening with distance. Another physical process at play is the energy-dependent diffusion of particles, in which high-energy particles diffuse more efficiently than the lower energy ones. An energy-dependent morphology can also be attributed to the bulk motion of particles in the PWN outflow. Advection-dominated transport in this case leads to a spatial distribution that is dependent on the energy loss timescale. Conversely, for SNRs with uniform gas densities, we expect the gamma-ray emission at lower energies to be concentrated to the central SNR position and the high-energy emission to be farther away (Celli et al., 2019; Brose et al., 2021). Relativistic jets, such as those observed in microquasars, are also expected to exhibit an energy-dependent morphology along their extent (H. E. S. S. Collaboration et al., 2024).
@@
The High Energy Stereoscopic System (H.E.S.S.) has detected several sources exhibiting energy-dependent morphologies. Examples include evolved PWNe such as HESS J1303--631 (H. E. S. S. Collaboration et al., 2012a) and HESS J1825--137, where the nebula size increases with decreasing energy. HESS J1825--137 is a well-known example of a PWN that shows a clear energy-dependent gamma-ray morphology (Aharonian et al., 2006a; H.E.S.S. Collaboration et al., 2019; Principe et al., 2020). A compact nebula with an extended diffuse nebula and an asymmetric morphology was discovered in X-rays towards PSR B1823--13 (Finley et al., 1996). Aharonian et al. (2005) discovered an extended very high-energy (VHE) gamma-ray nebula towards this region, designated HESS J1825--137. The size of the nebula is seen to increase with decreasing energy (H.E.S.S. Collaboration et al., 2019). At low energies (<32<32< 32 TeV) the emission nebula is extended, with its peak offset from the position of PSR B1823--13. For higher energies (>32>32> 32 TeV), the emission is more compact and close to the pulsar position. This behaviour indicates that the gamma-ray emission towards HESS J1825--137 originates from PSR B1823--13 as a PWN.
@@
SS 433 was the first microquasar to be detected at TeV energies (Abeysekara et al., 2018; H. E. S. S. Collaboration et al., 2024). The H.E.S.S. gamma-ray image, with the nearby extended emission from HESS J1908$+$063 modelled out, reveals two regions of emission, known as the eastern and western jets. No TeV emission is significantly detected from the central source. At the highest energies (¿10 TeV), the emission appears only in the base of the X-ray jets, i.e. the point closest to the central source. In contrast, the lower energy gamma rays (0.8–2.5 TeV and 2.5–10 TeV) peak in significance at positions farther along the jets.
@@
Various techniques, applicable in different scenarios, have been utilised to search for energy dependence. These include the comparison of spectral indices across different regions or energy bins in a map (Aharonian et al., 2006a), slicing spatial maps into different energy bands and comparing their extension to a model fit (see examples of Gaussian spatial model fits in H. E. S. S. Collaboration et al. 2012a, Principe et al. 2020, and MAGIC Collaboration et al. 2020), comparing radial profiles (Principe et al., 2020), and many other techniques. Energy-dependent morphology is a key observable for studying the nature of a VHE gamma-ray source. Here, we present a new tool in the Gammapy framework (Donath et al., 2023; Acero et al., 2024) for investigating the potential energy-dependent morphology towards a gamma-ray source. This method utilises maps across multiple energy bands to fit the morphological parameters of the source, such as position and extension. The user defines the best-fit model for the morphology and spectrum. A global fitting approach is applied, and then an individual fit is performed in each energy band to assess if the morphology changes significantly.
@@ 2 Methodology
In this section we describe various methods for measuring the energy dependence of gamma-ray sources, highlighting their respective advantages and disadvantages.
@@
One common approach is spatially resolved spectral analysis (Aharonian et al., 2006b, c). This technique can be used to search for a softening of the gamma-ray emission away from the pulsar position in a PWN scenario, which provides the evidence of an energy-dependent morphology. The signal region is divided into non-overlapping subregions. A spectral model, typically a power law, is fit in each subregion independently to search for changing photon index across the source region (Aharonian et al., 2006c). This approach offers a simple method of analysing the spectra, particularly for variations in spectral shape. However, this method has its limitations. The choice of region size and number of regions can strongly influence the result, leading to potential biases. The astrophysical background is difficult to determine accurately for each region, the estimation for different regions are typically not independent. These limitations are more significant for small subregions that approach the resolution of the point spread function (PSF), making this method better suited to more extended sources. Quantifying the significance of spectral variability across these subregions remains a challenge.
@@
Another method involves creating radial excess profiles along a specific axis (H. E. S. S. Collaboration et al., 2012b; H.E.S.S. Collaboration et al., 2019; Principe et al., 2020). The distribution of the emission is analysed by extracting the excess counts from specific regions in a specific orientation from the source position. This is useful to visualise changing spatial features such as extension and intensity. It is important to note that they do not take into account the PSF or variations in exposure. Radial profiles offer a model-independent method for assessing the spatial variations, allowing for an unbiased way of visualising the data. However, they assume radial symmetry within the region, which may not always be a valid approximation. Additionally, this method is highly sensitive to the selected axis direction of the profile and the choice of the radial bins. Another limitation is the difficulty in handling background systematics. Consequently, these dependencies make it a less robust analysis. Additionally, variations in the exposure across the field of view can change the excess distribution, meaning the excess partially reflects these exposure variations rather than the source properties itself. For example, a decrease in excess counts may be misinterpreted as a physical or temporal effect, when it is, in fact, stemming from the exposure variation. This makes deriving the true significance of the effect a challenge. While traditionally overlooked, accounting for exposure variation is now possible with the introduction of methods like the FluxProfileEstimator in Gammapy. This, however, still does not account for contributions from additional sources within the field of view or analysing variations on small spatial scales.
@@
We propose a new method that addresses each of the limitations described above, to quantify the energy-dependent morphology with a proper statistical assessment. This technique utilises a full forward-folding approach, where a morphological fit is conducted across various energy slices and compared with a global morphology fit. The technique chosen here involves fitting a spatial model and comparing the differences in morphological parameters (such as extension and position) in different energy bins. Further information about the tool and examples can be found in Sects. 3 and 4, respectively.
In gamma-ray astronomy we often utilise hypothesis testing. We rely on Wilks’ theorem (Wilks, 1938), which is a log-likelihood ratio test that allows us to estimate the relative significance through the test statistic ($\Delta\rm{TS}$),
@@
where $\mathcal{L}_{0}^{\rm{max}}$ and $\mathcal{L}_{1}^{\rm{max}}$ represent the maximum likelihood of the model under the null and alternative hypotheses, respectively. Under the correct conditions, where the two models are nested, and the regularity conditions are satisfied, we expect the distribution of $\Delta\rm{TS}$ to asymptotically approach a $\chi^{2}$ function with $m$ degrees of freedom, the difference in number of free parameters between the two nested hypotheses. Therefore, for a large enough sample it is possible to compute the $\Delta\rm{TS}$ and compare to a chi-squared function with the correct degrees of freedom to test the statistical significance.
For our tool, the null hypothesis (H0) is defined by a global fit. Initially, the morphological parameters are fit in the first energy band. The resulting parameters are then applied to all subsequent energy bands and a fit is performed across the different energy bands. The alternative hypothesis (H1) is defined by individual fits. Here, the dataset is divided into energy bins, and both the spatial and spectral models are fitted independently in each energy band. We apply Wilks’ theorem to these hypotheses to obtain the $\Delta\rm{TS}$ (see Eq. 1). The $\Delta\rm{TS}$ is converted into a significance value through the survival function of the chi-squared distribution. Thus, this method provides a quantitative measure of how much the morphology deviates from the null model for each energy band.
@@ 3 Implementation
To test for energy-dependent morphology using our tool, one begins by dividing the MapDataset into different energy bins, as specified by the user. Firstly, we allow (at a minimum) the position and extension of the SpatialModel to be free, along with the amplitude of the SpectralModel and normalisation of the background. It is up to the user to define which spatial parameters are to be kept free. For example, for a GaussianSpatialModel there are five free parameters — longitude, latitude, sigma, phi, and e. It is up to the user to make a judicious choice to freeze or free these parameters. In this case, longitude, latitude, and sigma are kept free. For a proper energy-dependent investigation, it is necessary to keep at least the position and extension free. This enables us to detect any change in the position or size of the gamma-ray source across different energy bands. The spectral index can be poorly constrained, particularly when fitting in small energy bands. Therefore, it is recommended to assume a given spectral shape while allowing the amplitude to be fit. This re-normalisation can account for any small variations in the spectrum.
Whilst there are several SpatialModels available in Gammapy, it is necessary to select a model that accurately represents the morphology of the gamma-ray source. SpatialModels such as the DiskSpatialModel and ShellSpatialModel, assume sharp edges, which are often not a good representation of the data, particularly across multiple energy bands. In contrast, the GaussianSpatialModel is able to represent the smooth features observed in gamma-ray morphology, particularly when the morphology varies with energy. This makes it a robust default for many analyses. Alternative models should be used with caution and only when clearly justified by the data. A usage example is shown in Fig. 1, corresponding to the example in Sect. 4.2.
@@ 4.1 Example 1: HESS J1825--137
To demonstrate the functionality of the tool, we showcase two examples. In the first example, we utilise a subset of the H.E.S.S. data towards HESS J1825--137 from Aharonian et al. (2006b). Three energy bins are chosen for this study: 0.4-2 TeV, 2-10 TeV and 10-100 TeV. For this specific example, a mask is applied to exclude the emission from the microquasar LS 5039 and the source HESS J1826--130. As shown by Fig. 2, the dataset shows three distinct energy bins.
@@
Figure 2: Significance map of the H.E.S.S. dataset towards HESS J1825--137 in the various energy bins of interest. An oversampling radius of $0.1^{\circ}$ is used. The white circles show the mask applied to the emission microquasar LS 5039 and the gamma-ray source HESS J1826--130. The turquoise dot indicates the position of the pulsar, PSR B1823--13. There is a clear change in the size of the emission in the different bins.
@@
The PowerLawSpectralModel and GaussianSpatialModel provide reasonable initial spectral and spatial models, respectively, for the dataset, so we used them for our study. The power law spectral shape is defined as follows:
$\phi(E)=\phi_{0}\cdot\left(\frac{E}{E_{0}}\right)^{-\Gamma},$ (2)
@@
where $\phi_{0}$ is the amplitude, $\Gamma$ is the index of the power law fit, and $E_{0}$ is a normalisation energy taken here to be 1 TeV. We utilise the GaussianSpatialModel in Gammapy, keeping the asymmetry frozen to 0 for simplicity, in the different energy bands, to test for any energy-dependent morphology with our tool.
@@
An important aspect of this new tool is to check that the energy bands chosen contain significant data. To do this, we can look at the signal above the background in each energy bin. This done by slicing the dataset into its energy bins. The null hypothesis (H0) involves freezing all parameters except for the normalisation of the background, which is fitted in each energy band. The alternative hypothesis (H1) allows additional parameters to be free, in this case the position and extension of the spatial model and the amplitude of the spectral model. Subsequently, the fit is performed for each energy band. These hypotheses are compared using Wilks’ theorem (Eq. 1) to obtain a $\Delta\rm{TS}$ that indicates the significance of the signal relative to the background in each energy band. If each energy bin is significant above the background ($\gg 5\sigma$), the energy-dependent study can continue.
@@
The next step is to quantify any energy-dependent morphology. The null hypothesis is determined by applying the same set of parameters to all energy bands and performing the fit. In the alternative hypothesis, the free parameters of the model are fit individually within each energy band. We plotted the fitted spatial models from our energy-dependent study. These are shown for each energy band in Fig. 3.
@@
Figure 3: Results of the estimator tool for the HESS J1825--137 dataset. The different coloured rings show the fitted GaussianSpatialModel in the various energy bins along with their uncertainties. The plus symbols show the position and error from the fit. The black dot indicates the position of the pulsar, PSR B1823--13.
@@
Aharonian et al. (2006b) investigated the spectral properties of HESS J1825--137 in different regions to search for any energy-dependent morphology. A fit to these results gives a chi-squared value of 58 with 11 degrees of freedom. This shows that the method outlined in our paper is more sensitive.